OpenStudy (anonymous):

Find the specified vector or scalar. u = -4i + 1j and v = 4i + 1j; Find ||u+v|| A. B. 8 C. 5 D. 2

OpenStudy (anonymous):

A. is square root of 34

OpenStudy (solomonzelman):

First, can you right the vector(u+v) ?

OpenStudy (anonymous):

im not sure what you mean

OpenStudy (solomonzelman):

-4i+4i=0i 1j+1j+2j So, vector(u+v) = 0i + 2j

OpenStudy (solomonzelman):

And then the magnitude of the vector u+v is: \(||u+v||=\sqrt{0^2+2^2}=?\)

OpenStudy (anonymous):

which would be 2?

OpenStudy (solomonzelman):

Yes, the magnitude of u+v is 2.

OpenStudy (solomonzelman):

(And this is scalar)

OpenStudy (anonymous):

thank you!

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

can you help me with another one? Ill start a thread and tag you